# Avoiding Catastrophe in Online Learning by Asking for Help

### v3

@@ A.2 Proof
Lemma 5.4 is a restatement and simplification of Lemma 3.5 in [33]. First, [33] parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$. Second, [33] include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound. In our case, an expected query bound is sufficient, so we simply set $k=\hat{k}$. Third, [33] provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it. Fourth, their number of actions $n$ is equal to $|\tilde{\Pi}|$ in our setting. (Their actions correspond to policies in $\tilde{\Pi}$, not our actions in $A$.) Since [33] set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$. Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
@@ B.2 Main proof
    Let $V_{T}=\{t\in[T]:\pi_{t}(s_{t})\neq\pi^{m}(s_{t})\}$ be the set of time steps where Hedge’s proposed action doesn’t match the mentor’s. Note that $|V_{T}|$ upper bounds the number of mistakes the algorithm makes (the number of mistakes could be smaller, since the algorithm sometimes queries instead of taking action $\pi_{t}(s_{t})$).
@@
    For $K\subseteq S$, let $\operatorname*{vol}(K)$ denote the $n$-dimensional Lebesgue measure of $K$.
@@
    We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\mathbf{s}$ and “Case 2” to refer to finite Littlestone dimension. In Case 1, expectations are over the randomness of both $\mathbf{s}$ and the algorithm, while in Case 2, expectations are over just the randomness of the algorithm. When we need to distinction, we use $\operatorname*{\mathbb{E}}_{\mathbf{a}}$ to denote the expectation over randomness of the algorithm and $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}$ to denote the expectation over $\mathbf{s}$.
@@
$\operatorname*{\mathbb{E}}[|V_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
@@
Define $\ell:[T]\times\tilde{\Pi}\to[0,1]$ by $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$, and let $w^{h}$ and $\pi_{t}^{h}$ denote the values of $w$ and $\pi_{t}$ respectively in HedgeWithQueries, while $w$ and $\pi_{t}$ refer to the variables in Algorithm 1. Then $w$ and $w^{h}$ evolve in the exact same way, so the distributions of $\pi_{t}$ and $\pi_{t}^{h}$ coincide. Thus by Lemma 5.4,
$\operatorname*{\mathbb{E}}_{\mathbf{a}}\left[\sum_{t=1}^{T}\ell(t,\pi_{t})\right]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\tilde{\pi})\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}=2T\varepsilon\log|\tilde{\Pi}|$
Since Lemma 5.4 holds for any loss function, the bound above holds for any $\mathbf{s}\in S^{T}$, so the bound also holds in expectation over $\mathbf{s}\sim\boldsymbol{\mathcal{D}}$ (which is needed for Case 1). Next, observe that $|V_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(s_{t})\neq\pi^{m}(s_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, so
$\operatorname*{\mathbb{E}}_{\mathbf{a}}[|V_{T}|]\leq 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))\right]\leq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\sup_{\pi\in\Pi}\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi(s_{t}))\right]$
@@
with the first step due to $\pi^{m}\in\Pi$ and the second step due to Lemma 5.3. The last component we need is that $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.1). Combining the above inequalities and taking the expectation over $\mathbf{s}\sim\boldsymbol{\mathcal{D}}$ (in addition to the randomness of the algorithm), we get
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}},\,\mathbf{a}}\,[|V_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))\right]$
@@
Case 2: Since $\tilde{\Pi}$ is an adversarial cover of $\Pi$ and $\pi^{m}\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))=0$. Since $|\tilde{\Pi}|\leq(eT/d)^{d}$ (with such a $\tilde{\Pi}$ guaranteed to exist by Lemma 5.2),
$\operatorname*{\mathbb{E}}_{\mathbf{a}}[|V_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))$
@@
Fix any $t\in[T]$. If $t\in Q_{T}$, then $\mu(s_{t},a_{t})=\mu^{m}(s_{t})$, so assume $t\not\in Q_{T}$. Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}:\pi_{t}(s_{t})=a}||s_{t}-s||$. Since $t\not\in Q_{T}$, we must have $||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$.
We have $a^{\prime}=\pi^{m}(s^{\prime})$ by construction of $X_{t}$ and $\pi_{t}(s_{t})=a^{\prime}$ by construction of $a^{\prime}$. Combining these with the local generalization assumption, we get
$\mu(s_{t},a_{t})=\mu(s_{t},\pi_{t}(s_{t}))=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$
@@
Assume $x_{1},\dots,x_{T},y_{1},\dots,y_{T}\in[0,1]$ and $x_{t}\geq y_{t}$ for all $t\in[T]$. Then
$\prod_{t=1}^{T}x_{t}-\prod_{t=1}^{T}y_{t}\leq\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}$
@@
We proceed by induction on $T$. The claim is trivially satisfied for $T=1$, so suppose $T>1$ and assume that $\prod_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}y_{t}\leq\sum_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}$. Then
$\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}-\prod_{t=1}^{T}x_{t}+\prod_{t=1}^{T}y_{t}=$ $\ x_{T}\sum_{t=1}^{T-1}x_{t}-y_{T}\sum_{t=1}^{T-1}y_{t}-x_{T}\prod_{t=1}^{T-1}x_{t}+y_{T}\prod_{t=1}^{T-1}y_{t}$
$=$ $\ x_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
Since $T>1$ and $x_{t}\in[0,1]$ for all $t\in[T]$, we have $\sum_{t=1}^{T-1}x_{t}\geq x_{1}\geq\sum_{t=1}^{T-1}x_{t}$. Thus $\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\geq 0$. Combining this with $x_{T}\geq y_{T}$, we get
$\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}-\prod_{t=1}^{T}x_{t}+\prod_{t=1}^{T}y_{t}=$ $\ x_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
$\geq$ $\ y_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
$=$ $\ y_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}+\prod_{t=1}^{T-1}y_{t}\right)$
@@
The last step is due to $y_{T}\geq 0$ and our assumption of $\prod_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}y_{t}\leq\sum_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}$. ∎
@@
We first claim that $a_{t}=\pi^{m}(s_{t})$ for all $t\not\in V_{T}$. If $t\in Q_{T}$, the claim is immediate; if not, we have $a_{t}=\pi_{t}(s_{t})$, and $\pi_{t}(s_{t})=\pi^{m}(s_{t})$ due to $t\not\in V_{T}$. Thus $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=\mu^{m}(s_{t})$ for $t\not\in V_{T}$.
We next claim that $\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq L\varepsilon^{1/n}$ for all $t\in[T]$. If $\mu(s_{t},a_{t})\leq\mu^{m}(s_{t})$, this follows from Lemma B.2. If $\mu(s_{t},a_{t})>\mu^{m}(s_{t})$, then $\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=0\leq L\varepsilon^{1/n}$. Therefore
$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq$ $\ \sum_{t\in V_{T}}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
$\leq$ $\ \sum_{t\in V_{T}}L\varepsilon^{1/n}$
$=$ $\ |V_{T}|L\varepsilon^{1/n}$
Now let $x_{t}=\mu^{m}(s_{t})$ and $y_{t}=\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
$\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
Since $\mu(s_{t},a_{t})\geq\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$, we have
$R_{T}=$ $\ \prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
$\leq$ $\ |V_{T}|L\varepsilon^{1/n}$
Taking the expectation and applying Lemma B.1 completes the proof. ∎
@@
Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$. Then $W\subseteq K$ is a $\delta$-packing of $K$ if for all $x,y\in W$, $||x-y||>\delta$. The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
@@
Lemma B.6 (Jung’s Theorem [22]).
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$

### v4

@@ A.2 Proof
Lemma 5.4 is a restatement and simplification of Lemma 3.5 in Russo et al. (2024). First, Russo et al. (2024) parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$. Second, Russo et al. (2024) include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound. In our case, an expected query bound is sufficient, so we simply set $k=\hat{k}$. Third, Russo et al. (2024) provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it. Fourth, their number of actions $n$ is equal to $|\tilde{\Pi}|$ in our setting. (Their actions correspond to policies in $\tilde{\Pi}$, not our actions in $\mathcal{Y}$.) Since Russo et al. (2024) set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$. Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
@@ B.2 Main proof
    Let $M_{T}=\{t\in[T]:\pi_{t}(x_{t})\neq\pi^{m}(x_{t})\}$ be the set of time steps where Hedge’s proposed action doesn’t match the mentor’s. Note that $|M_{T}|$ upper bounds the number of mistakes the algorithm makes (the number of mistakes could be smaller, since the algorithm sometimes queries instead of taking action $\pi_{t}(x_{t})$).
@@
    For $X\subseteq\mathcal{X}$, let $\operatorname*{vol}(X)$ denote the $n$-dimensional Lebesgue measure of $X$.
@@
    We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\boldsymbol{x}$ and “Case 2” to refer to finite Littlestone dimension. In Case 1, expectations are over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, expectations are over just the randomness of the algorithm. When we need to distinguish, we use $\operatorname*{\mathbb{E}}_{\boldsymbol{y}}$ to denote the expectation over randomness of the algorithm and $\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}$ to denote the expectation over $\boldsymbol{x}$.
@@
$\operatorname*{\mathbb{E}}[|M_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
@@
Define $\ell:[T]\times\tilde{\Pi}\to[0,1]$ by $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$, and let $w^{h}$ and $\pi_{t}^{h}$ denote the values of $w$ and $\pi_{t}$ respectively in HedgeWithQueries, while $w$ and $\pi_{t}$ refer to the variables in Algorithm 1. Then $w$ and $w^{h}$ evolve in the exact same way, so the distributions of $\pi_{t}$ and $\pi_{t}^{h}$ coincide. Thus by Lemma 5.4,
$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}\left[\sum_{t=1}^{T}\ell(t,\pi_{t})\right]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\tilde{\pi})\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}=2T\varepsilon\log|\tilde{\Pi}|$
Since Lemma 5.4 holds for any loss function, the bound above holds for any $\boldsymbol{x}\in S^{T}$, so the bound also holds in expectation over $\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}$ (which is needed for Case 1). Next, observe that $|M_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(x_{t})\neq\pi^{m}(x_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, so
$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}[|M_{T}|]\leq 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
@@
$\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))\right]\leq$ $\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\sup_{\pi\in\Pi}\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi(x_{t}))\right]$
@@
with the first step due to $\pi^{m}\in\Pi$ and the second step due to Lemma 5.3. The last component we need is that $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.1). Combining the above inequalities and taking the expectation over $\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}$ (in addition to the randomness of the algorithm), we get
$\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}},\,\boldsymbol{y}}\,[|M_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))\right]$
@@
Case 2: Since $\tilde{\Pi}$ is an adversarial cover of $\Pi$ and $\pi^{m}\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))=0$. Since $|\tilde{\Pi}|\leq(eT/d)^{d}$ (with such a $\tilde{\Pi}$ guaranteed to exist by Lemma 5.2),
$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}[|M_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
@@
Consider an arbitrary $t\in[T]$. If $t\in Q_{T}$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$, so assume $t\not\in Q_{T}$. Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in X_{t}:\pi_{t}(x_{t})=y}||x_{t}-x||$. Since $t\not\in Q_{T}$, we must have $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$.
We have $y^{\prime}=\pi^{m}(x^{\prime})$ by construction of $X_{t}$ and $\pi_{t}(x_{t})=y^{\prime}$ by construction of $y^{\prime}$. Combining these with the local generalization assumption, we get
$\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi_{t}(x_{t}))=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
@@
Assume $a_{1},\dots,a_{T},b_{1},\dots,b_{T}\in[0,1]$ and $a_{t}\geq b_{t}$ for all $t\in[T]$. Then
$\prod_{t=1}^{T}a_{t}-\prod_{t=1}^{T}b_{t}\leq\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}$
@@
We proceed by induction on $T$. The claim is trivially satisfied for $T=1$, so suppose $T>1$ and assume that $\prod_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}b_{t}\leq\sum_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}$. Then
$\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}-\prod_{t=1}^{T}a_{t}+\prod_{t=1}^{T}b_{t}=$ $\ a_{T}\sum_{t=1}^{T-1}a_{t}-b_{T}\sum_{t=1}^{T-1}b_{t}-a_{T}\prod_{t=1}^{T-1}a_{t}+b_{T}\prod_{t=1}^{T-1}b_{t}$
$=$ $\ a_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
Since $T>1$ and $a_{t}\in[0,1]$ for all $t\in[T]$, we have $\sum_{t=1}^{T-1}a_{t}\geq a_{1}\geq\sum_{t=1}^{T-1}a_{t}$. Thus $\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\geq 0$. Combining this with $a_{T}\geq b_{T}$, we get
$\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}-\prod_{t=1}^{T}a_{t}+\prod_{t=1}^{T}b_{t}=$ $\ a_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
$\geq$ $\ b_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
$=$ $\ b_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}+\prod_{t=1}^{T-1}b_{t}\right)$
@@
The last step is due to $b_{T}\geq 0$ and our assumption of $\prod_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}b_{t}\leq\sum_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}$. ∎
@@
$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
@@
We first claim that $y_{t}=\pi^{m}(x_{t})$ for all $t\not\in M_{T}$. If $t\in Q_{T}$, the claim is immediate; if not, we have $y_{t}=\pi_{t}(x_{t})$, and $\pi_{t}(x_{t})=\pi^{m}(x_{t})$ due to $t\not\in M_{T}$. Thus $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=\mu_{t}^{m}(x_{t})$ for $t\not\in M_{T}$.
We next claim that $\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq L\varepsilon^{1/n}$ for all $t\in[T]$. If $\mu_{t}(x_{t},y_{t})\leq\mu_{t}^{m}(x_{t})$, this follows from Lemma B.2. If $\mu_{t}(x_{t},y_{t})>\mu_{t}^{m}(x_{t})$, then $\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=0\leq L\varepsilon^{1/n}$. Therefore
$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq$ $\ \sum_{t\in M_{T}}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
$\leq$ $\ \sum_{t\in M_{T}}L\varepsilon^{1/n}$
$=$ $\ |M_{T}|L\varepsilon^{1/n}$
Now let $a_{t}=\mu_{t}^{m}(x_{t})$ and $b_{t}=\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
$\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
Since $\mu_{t}(x_{t},y_{t})\geq\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$, we have
$R_{T}=$ $\ \prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
$\leq$ $\ |M_{T}|L\varepsilon^{1/n}$
Since we also have $\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}(\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t})))$, we get both of the following
$R_{T}\leq$ $\ L\varepsilon^{1/n}|M_{T}|$
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ L\varepsilon^{1/n}|M_{T}|$
Taking the expectation and applying Lemma B.1 to both of the above inequalities completes the proof. ∎
@@
Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$. Then $X\subseteq K$ is a $\delta$-packing of $K$ if for all $x,y\in X$, $||x-y||>\delta$. The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
@@
Lemma B.6 (Jung’s Theorem (Jung, 1901)).
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$